# Exploring Exponential Runge-Kutta Methods: A Survey

### v1

@@ 1. Introduction
where $u(t)\in\mathbb{R}^{d}$, $t\geq 0$, $A$ is a linear operator (typically a matrix) describing the (typically stiff) part of the problem, while $g$ represents the (typically non-stiff) nonlinear component. Remarkably, despite its fundamental importance, there is no univocal and universally accepted definition of stiffness. Some mathematicians characterize it through the presence of components of the solution that evolve on very different time scales, others define it in terms of the ratio between the eigenvalues of the Jacobian matrix of the system, while still others relate it to the practical difficulties that arise when using explicit numerical methods. This multiplicity of definitions reflects the intrinsic complexity of the phenomenon and the different perspectives from which it can be analyzed. We shall provide a more exhaustive discussion about the topic in Sect. 2.1.
@@
Runge-Kutta (RK) methods represent a milestone in the development of numerical methods for ordinary differential equations. Initially introduced by Carl Runge in 1895 [80] and later refined by Martin Wilhelm Kutta in 1901 [54], these methods revolutionized the approach to the numerical solution of ODEs. Their classical formulation, based on a linear combination of function evaluations at appropriately chosen intermediate points, offers a remarkable balance between accuracy and simplicity of implementation. However, despite their historical success, classical explicit RK methods have shown significant limitations when applied to more complex or stiff problems. [^1] It is important here to make a necessary distinction between explicit and implicit RK methods. Explicit methods compute each stage using only previously computed ones, but suffer from strong stability limitations, especially in the presence of stiff problems. On the contrary, implicit methods guarantee better stability and are particularly suitable for tackling stiff problems. The systematic introduction of implicit methods dates back to the pioneering works of John Butcher [11], who provided a solid theoretical basis for their development and application. The main drawback of implicit methods is their high computational cost, since they require the knowledge of the solution for nonlinear systems of equations at each stage. The need to use extremely small time steps to ensure stability in the already mentioned stiff problems, with the consequent increase in computational cost, has pushed the scientific community towards the development of more sophisticated variants. Furthermore, the inability of classical RK methods to naturally preserve some important physical properties of the systems (such as invariants and energies) has motivated the search for more advanced geometric and structural methods. [^2] We remark that in the mid-1990s, two important reviews regarding RK methods were published. The first one, dated 1994, is due to Enright et al. [30] and it focuses on algorithms related to explicit methods, while the second one was written by Butcher in 1996 [12] and with a more historical approach. Butcher concentrates on the evolution of the methods from their origins, with particular attention to implicit methods. Another interesting survey about Pseudo-RK methods is due to Costabile [18] and it was published in 2003. We remark that in none of these works is there any reference to exponential methods.
@@
Thus, the presence of stiff terms represents a fascinating challenge for traditional numerical methods. Among the various approaches developed to address this class of problems, exponential RK (ExpRK) methods have assumed a particularly important role. As pointed out by Hochbruck, Lubich and Selhofer [43]:
@@
The efficient implementation of these methods requires two fundamental ingredients: the computation of $\mathrm{e}^{A}$ and the approximation of the integral containing the nonlinear term. The matrix exponential can be characterized and computed in different ways: through the classical power series, through spectral decomposition when possible, or through integral representations in the complex plane (see, for instance, [39]). The numerical evaluation of the matrix exponential is an active area of research itself, with techniques ranging from scaling and squaring methods to approaches based on Krylov subspaces, pioneered by Saad [81]. As for the integral term, its approximation is achieved through suitable combinations of evaluations of the function $g$, weighted through the so-called $\varphi$-functions. These functions, which naturally generalize the matrix exponential, emerge from the analysis of the integral term and play a central role in the construction and analysis of these methods. We highlight that $\varphi$-functions can be defined in several equivalent ways. A particularly useful definition is through the following integral:
@@
This article pursues two main aims. On one hand, we shall provide a comprehensive overview of ExpRK methods, tracing their historical development from their origins to the most recent theoretical developments in contemporary research. On the other hand, we propose teaching tools for those who approach these methods for the first time: starting from the fundamental theoretical bases, the article gradually develops to the most advanced aspects of the subject, supported by explanatory examples that illustrate their main characteristics.
The manuscript is organized as follows. Section 2 provides the complete theoretical framework of ExpRK. Section 3 presents a historical overview of the fundamental theoretical results, starting from the early works up to the most recent review of exponential methods by Hochbruck and Ostermann [46]. In Section 4, we recapitulate the most recent developments of exponential Runge-Kutta methods, from their applications to high stiff order and delay differential equations up to their usage in several modeling frameworks. Finally, Section 5 is devoted to the numerical simulations of two test cases of low dimension that illustrate the discussed theoretical properties.
@@ 2.1. Stiffness in a nutshell
Despite the word stiff having been used for several decades, defining stiffness in precise mathematical terms seems to be an impossible task. To the best of the authors’ knowledge, the term first appeared in [20], where the phenomenon of stiffness was shown by means of a scalar example. One could try to define stiffness in terms of the inherent characteristics of the problem at hand, without explicitly referring to its numerical approximation. From this perspective, two large (not mutually exclusive) groups of problems may be classified as stiff:
@@
Naturally, the above are not rigorous definitions, nor are they truly independent of the numerical methods considered. In particular, “large” in the first group is somehow related to the stepsize $h$: it would be more accurate to say that a problem is stiff if $|\lambda h|$ is large, or, in other words, if it requires “unreasonable restrictions on $h$[74]. As for the second group, observe that the fast and slow components are usually associated to the large and small (in modulus) eigenvalues of the Jacobian, respectively [22] and, indeed, the ratio between the largest and the smallest real parts - the stiffness ratio - was once considered a possible measure of stiffness [55].
While both characterizations above concern properties which are typical of stiff problems, it has become clear over time that neither represents an exhaustive explanation of the complex phenomenon. More pragmatically, it is now deemed more appropriate to define stiffness in terms of the numerical issues arising from the phenomenon. In other words [36]:
@@
Or, at the very least, problems for which explicit methods perform worse than implicit ones. For an insightful discussion on the topic of stiffness, we refer the interested readers to [56]. It is shown therein, by means of numerical experiments, that the stiffness ratio cannot be interpreted as a measure of stiffness and the classic pseudodefinitions are neither sufficient nor necessary conditions for stiffness. Moreover, the pragmatic definition is also rephrased so as to emphasize that stiffness varies with the specific interval over which the problem is integrated:
@@ 2.2. Constructing ExpRK methods
At the end of the 1980s, Gallopoulos and Saad [32, 33] were among the first to emphasize that in the context of numerical integration one does not need to actually compute exponentials of matrices, but rather only products of the form $\mathrm{e}^{-A}v$, where $v$ is a vector. This led them to explore the possibility to apply Krylov methods in this context. The approach consists in approximating $\mathrm{e}^{-A}v$ with a member of the Krylov subspace
@@
[…]
@@
Thus, the problem reduces to computing the exponential of a much smaller matrix than the original one. A different approach proposed in [32] consists in approximating the exponential with a rational function. As highlighted in [33], the two approaches can be combined, applying the rational approximation to $\mathrm{e}^{-H_{m}}$ while making use of (5). Eventually, Saad [81] further improved the accuracy obtained with the Krylov approach by rewriting $\mathrm{e}^{-A}v=v-A\varphi_{1}(-A)v$ and computing a Krylov approximation for $\varphi_{1}(-A)$ instead of one for $\mathrm{e}^{-A}$.
@@
Later, Beylkin, Keiser and Vozovoi [9] observed that the problem with Taylor truncation methods (so, also with Padé rational approximations, whose numerators and denominators are truncated Taylor series) arise when the matrix $A$ possesses large singular values. Thus, they proposed the scaling and squaring algorithm to evaluate the exponentials, which consists in exploiting the equality $\mathrm{e}^{A}=(\mathrm{e}^{\frac{A}{2^{s}}})^{2^{s}}$, where $s$ is chosen such that the largest singular value of $\mathrm{e}^{\frac{A}{2^{s}}}$ is less than one. The authors also made use of this approach to compute all $\varphi_{k}$ functions by means of their recursive formula (4). The scaling and squaring method is probably the most popular for computing matrix exponentials and it is used in MATLAB’s command expm in combination with a Padé approximation using the same degree for the numerator and denominator [38].
@@
In 2003, Davies and Higham [24] proposed yet a different approach, based on decomposing the matrix $A$ as $A=S^{-1}BS$ such that $\mathrm{e}^{B}$ (or $\varphi_{k}(B)$) is much easier to compute than $\mathrm{e}^{A}$, and $S$ has a reasonably low condition number. In particular, their idea was to use a Schur decomposition (such that $S$ is unitary and $B$ is upper triangular) combined with a recurrence algorithm of Parlett [78] which exploits the fact that functions of block upper triangular matrices inherit their block structure. If $A$ is symmetric, the approach introduced by Lu [58] for computing $\varphi_{1}(A)$ turns out to be more efficient. The method is again based on a matrix decomposition, obtained with an orthogonal $S$ and a tridiagonal $B$. The idea in this case is to approximate $\mathrm{e}^{B}$ using a Chebyshev rational approximation for $\mathrm{e}^{B-\lambda I}$, where $\lambda$ is the largest eigenvalue of $B$. The algorithm has then been generalized to compute all $\varphi$-functions [72].
@@
Another class of methods that can be employed to approximate the $\varphi$-functions is that of polynomial interpolation, such as interpolation on Leja points [15].
@@
A straightforward application of (4) to compute the $\varphi$-functions suffers from cancellation errors for small $z$. The idea implemented by Cox and Matthews [19] was to use a piecewise strategy, namely to compute the $\varphi$-functions directly with (4) for large $z$, and with truncated Taylor series for small $z$. There is a certain region, however, where neither approach is accurate enough. Kassam and Trefethen [51] propose a method with the idea of bridging the gap between the “small” and “large” $z$, based on the computation of the contour integral
@@
In 2009, Al-Mohy and Higham [1] proposed a heuristic improvement on the standard scaling and squaring algorithm combined with Padé approximation. The goal was to solve the problem of overscaling, namely the fact that for large $\|A\|$ a value of $s$ is chosen which is too large, causing loss of accuracy in floating-point arithmetic. The modification involves exploiting the values $\|A^{k}\|^{\frac{1}{k}}$ and the improvements are more evident for upper triangular matrices.
The (currently) most popular method to compute the $\varphi$-functions was proposed by the same authors in 2011 [2]. In fact, the purpose of that paper is twofold. The authors observe first that the problem of computing sums of the form $\sum_{k=0}^{p}\varphi_{k}(A)u_{k}$ functions defining an exponential method can be recasted as a single matrix exponential $e^{\tilde{A}}$ where $\tilde{A}$ is obtained by suitably augmenting $A$ with $p$ rows and columns. Next, they propose their algorithm for computing products of matrix exponential times matrices $\mathrm{e}^{A}B$ for $B\in\mathbb{R}^{n\times m}$ with $m\ll n$, as well as $\mathrm{e}^{tA}B$ where $t$ is an array of equally-spaced time points. It is based on the scaling part of the scaling and squaring, plus a truncated Taylor series.
@@
Optimal values for $s$ and $m$ are obtained through a backward error analysis. The algorithm incorporates the ideas previously proposed in [1] in order to minimize overscaling and turns out to be generally superior to existing ones in terms of accuracy, stability and predictability of cost, since the only convergence test involved is a test for early termination built into the evaluation of the truncated Taylor series. [^3] We remark that the algorithm is implemented in the MATLAB function expmv.
@@
$u_{1}$ $=(4.54\times 10^{-5})\times 1+0.01\varphi_{1}(-10)\times 1\approx 5.54\times 10^{-5}$
Now, we can compare the performance of ETD Euler and the classical fourth-order RK method (RK4), whose application to (6) gives:
@@
Observe that the origin is a stable steady state of a numerical method for the linear equation whenever $|h\lambda|$ lies in the corresponding stability region. Thus, in general, stability of the origin depends on the choice of $h$. A desirable property of a numerical method is $A$-stability: a scheme is $A$-stable [21] if the entire negative complex half plane is contained in its stability domain, such that the stability of the origin is independent of the chosen step size.

### v2

@@ 1. Introduction
where $u(t)\in\mathbb{R}^{d}$, $t\geq 0$, AA is a linear operator (typically a matrix) describing the (typically stiff) part of the problem, while $g$ represents the (typically non-stiff) nonlinear component. Remarkably, despite its fundamental importance, there is no univocal and universally accepted definition of stiffness. Some mathematicians characterize it through the presence of components of the solution that evolve on very different time scales, others define it in terms of the ratio between the eigenvalues of the Jacobian matrix of the system, while still others relate it to the practical difficulties that arise when using explicit numerical methods. This multiplicity of definitions reflects the intrinsic complexity of the phenomenon and the different perspectives from which it can be analyzed. We shall provide a more exhaustive discussion about the topic in Sect. 2.1.
@@
Runge-Kutta (RK) methods represent a milestone in the development of numerical methods for ordinary differential equations. Initially introduced by Carl Runge in 1895 [runge1895] and later refined by Martin Wilhelm Kutta in 1901 [kutta1901], these methods revolutionized the approach to the numerical solution of ODEs. Their classical formulation, based on a linear combination of function evaluations at appropriately chosen intermediate points, offers a remarkable balance between accuracy and simplicity of implementation. However, despite their historical success, classical explicit RK methods have shown significant limitations when applied to more complex or stiff problems. [^1] It is important here to make a necessary distinction between explicit and implicit RK methods. Explicit methods compute each stage using only previously computed ones, but suffer from strong stability limitations, especially in the presence of stiff problems. On the contrary, implicit methods guarantee better stability and are particularly suitable for tackling stiff problems. The systematic introduction of implicit methods dates back to the pioneering works of John Butcher [Butcher64], who provided a solid theoretical basis for their development and application. The main drawback of implicit methods is their high computational cost, since they require the knowledge of the solution for nonlinear systems of equations at each stage. The need to use extremely small time steps to ensure stability in the already mentioned stiff problems, with the consequent increase in computational cost, has pushed the scientific community towards the development of more sophisticated variants. Furthermore, the inability of classical RK methods to naturally preserve some important physical properties of the systems (such as invariants and energies) has motivated the search for more advanced geometric and structural methods. [^2] We remark that in the mid-1990s, two important reviews regarding RK methods were published. The first one, dated 1994, is due to Enright et al. [enright1994survey] and it focuses on algorithms related to explicit methods, while the second one was written by Butcher in 1996 [butcher1996history] and with a more historical approach. Butcher concentrates on the evolution of the methods from their origins, with particular attention to implicit methods. Another interesting survey about Pseudo-RK methods is due to Costabile [costabile2003survey] and it was published in 2003. We remark that in none of these works is there any reference to exponential methods.
@@
Thus, the presence of stiff terms represents a fascinating challenge for traditional numerical methods. Among the various approaches developed to address this class of problems, exponential RK (ExpRK) methods have assumed a particularly important role. As pointed out by Hochbruck, Lubich and Selhofer [hochbruck1998exponential]:
@@
The efficient implementation of these methods requires two fundamental ingredients: the computation of $\mathrm{e}^{A}$ and the approximation of the integral containing the nonlinear term. The matrix exponential can be characterized and computed in different ways: through the classical power series, through spectral decomposition when possible, or through integral representations in the complex plane (see, for instance, [higham2008functions]). The numerical evaluation of the matrix exponential is an active area of research itself, with techniques ranging from scaling and squaring methods to approaches based on Krylov subspaces, pioneered by Saad [saad1992analysis]. As for the integral term, its approximation is achieved through suitable combinations of evaluations of the function $g$, weighted through the so-called $\varphi$-functions. These functions, which naturally generalize the matrix exponential, emerge from the analysis of the integral term and play a central role in the construction and analysis of these methods. We highlight that $\varphi$-functions can be defined in several equivalent ways. A particularly useful definition is through the following integral:
@@
This article pursues a dual objective. First, it provides a comprehensive overview of exponential Runge–Kutta methods, tracing their development from the earliest formulations to some of the most recent theoretical advances. The goal is to offer a clear and coherent account of how these methods have evolved, why they remain of current interest, and what their key characteristics are, while keeping the presentation concise and focused on the essential points. Second, the article is intended as a resource for readers encountering these methods for the first time. Starting from the fundamental theoretical principles, the exposition gradually progresses to more advanced topics, accompanied by illustrative examples that clarify the main features and practical significance of the methods. By combining these two perspectives, the survey aims to be both a thorough reference for specialists and an accessible introduction for newcomers, providing a presentation that is at once comprehensive, structured, and approachable.
The manuscript is organized as follows. Section 2 provides the complete theoretical framework of ExpRK. Section 3 presents a historical overview of the fundamental theoretical results, starting from the early works up to the most recent review of exponential methods by Hochbruck and Ostermann [hochbruck2010exponential]. In Section 4, we recapitulate the most recent developments of exponential Runge-Kutta methods, from their applications to high stiff order and delay differential equations up to their usage in several modeling frameworks. Finally, Section 5 is devoted to the numerical simulations of three test cases, two of which of low dimension, that illustrate the discussed theoretical properties.
@@ 2.1. Stiffness in a nutshell
Despite the word stiff having been used for several decades, defining stiffness in precise mathematical terms seems to be an impossible task. To the best of the authors’ knowledge, the term first appeared in [curtiss1952integration], where the phenomenon of stiffness was shown by means of a scalar example. One could try to define stiffness in terms of the inherent characteristics of the problem at hand, without explicitly referring to its numerical approximation. From this perspective, two large (not mutually exclusive) groups of problems may be classified as stiff:
@@
Naturally, the above are not rigorous definitions, nor are they truly independent of the numerical methods considered. In particular, “large” in the first group is somehow related to the step size $h$: it would be more accurate to say that a problem is stiff if $|\lambda h|$ is large, or, in other words, if it requires “unreasonable restrictions on hh[miranker1981stiffness]. As for the second group, observe that the fast and slow components are usually associated to the large and small (in modulus) eigenvalues of the Jacobian, respectively [dahlquist1973stiffness] and, indeed, the ratio between the largest and the smallest real parts - the stiffness ratio - was once considered a possible measure of stiffness [lambert1973stiffratio].
While both characterizations above concern properties which are typical of stiff problems, it has become clear over time that neither represents an exhaustive explanation of the complex phenomenon. More pragmatically, it is now deemed more appropriate to define stiffness in terms of the numerical issues arising from the phenomenon. In other words [hw91]:
@@
Or, at the very least, problems for which explicit methods perform worse than implicit ones. For an insightful discussion on the topic of stiffness, we refer the interested readers to [lambert1991numerical]. It is shown therein, by means of numerical experiments, that the stiffness ratio cannot be interpreted as a measure of stiffness and the classic pseudodefinitions are neither sufficient nor necessary conditions for stiffness. Moreover, the pragmatic definition is also rephrased so as to emphasize that stiffness varies with the specific interval over which the problem is integrated:
@@ 2.2. Constructing ExpRK methods
At the end of the 1980s, Gallopoulos and Saad [gallopoulos1989parallel, gallopoulos1992efficient] were among the first to emphasize that in the context of numerical integration one does not need to actually compute exponentials of matrices, but rather only products of the form $\mathrm{e}^{-A}v$, where $v$ is a vector. This led them to explore the possibility to apply Krylov methods in this context. The approach consists in approximating $\mathrm{e}^{-A}v$ with a member of the Krylov subspace
@@
(5) $\mathrm{e}^{-A}v\approx V_{m}\mathrm{e}^{-H_{m}}\mathbf{1}_{m},\qquad\mathbf{1}_{m}=[1\,1\,\cdots 1]^{\top}\in\mathbb{R}^{m}.$
@@
Thus, the problem reduces to computing the exponential of a much smaller matrix than the original one. A different approach proposed in [gallopoulos1989parallel] consists in approximating the exponential with a rational function. As highlighted in [gallopoulos1992efficient], the two approaches can be combined, applying the rational approximation to $\mathrm{e}^{-H_{m}}$ while making use of (5). Eventually, Saad [saad1992analysis] further improved the accuracy obtained with the Krylov approach by rewriting $\mathrm{e}^{-A}v=v-A\varphi_{1}(-A)v$ and computing a Krylov approximation for $\varphi_{1}(-A)$ instead of one for $\mathrm{e}^{-A}$.
@@
Later, Beylkin, Keiser and Vozovoi [beylkin1998new] observed that the problem with Taylor truncation methods (so, also with Padé rational approximations, whose numerators and denominators are truncated Taylor series) arise when the matrix $A$ possesses large singular values. Thus, they proposed the scaling and squaring algorithm to evaluate the exponentials, which consists in exploiting the equality $\mathrm{e}^{A}=(\mathrm{e}^{\frac{A}{2^{s}}})^{2^{s}}$, where $s$ is chosen such that the largest singular value of $\mathrm{e}^{\frac{A}{2^{s}}}$ is less than one. The authors also made use of this approach to compute all $\varphi_{k}$ functions by means of their recursive formula (4). The scaling and squaring method is probably the most popular for computing matrix exponentials and it is used in MATLAB’s command expm in combination with a Padé approximation using the same degree for the numerator and denominator [higham2005scaling].
@@
In 2003, Davies and Higham [davies2003schur] proposed yet a different approach, based on decomposing the matrix $A$ as $A=S^{-1}BS$ such that $\mathrm{e}^{B}$ (or $\varphi_{k}(B)$) is much easier to compute than $\mathrm{e}^{A}$, and $S$ has a reasonably low condition number. In particular, their idea was to use a Schur decomposition (such that $S$ is unitary and $B$ is upper triangular) combined with a recurrence algorithm of Parlett [parlett1974computation] which exploits the fact that functions of block upper triangular matrices inherit their block structure. If AA is symmetric, the approach introduced by Lu [lu2003computing] for computing $\varphi_{1}(A)$ turns out to be more efficient. The method is again based on a matrix decomposition, obtained with an orthogonal $S$ and a tridiagonal $B$. The idea in this case is to approximate $\mathrm{e}^{B}$ using a Chebyshev rational approximation for $\mathrm{e}^{B-\lambda I}$, where $\lambda$ is the largest eigenvalue of $B$. The algorithm has then been generalized to compute all $\varphi$-functions [minchevthesis].
@@
Another class of methods that can be employed to approximate the $\varphi$-functions is that of polynomial interpolation, such as interpolation on Leja points [CALIARI200479].
@@
A straightforward application of (4) to compute the $\varphi$-functions suffers from cancellation errors for small $z$. The idea implemented by Cox and Matthews [cox2002exponential] was to use a piecewise strategy, namely to compute the $\varphi$-functions directly with (4) for large $z$, and with truncated Taylor series for small $z$. There is a certain region, however, where neither approach is accurate enough. Kassam and Trefethen [kassam2005fourth] propose a method with the idea of bridging the gap between the “small” and “large” $z$, based on the computation of the contour integral
@@
In 2009, Al-Mohy and Higham [al2010new] proposed a heuristic improvement on the standard scaling and squaring algorithm combined with Padé approximation. The goal was to solve the problem of overscaling, namely the fact that for large $\|A\|$ a value of $s$ is chosen which is too large, causing loss of accuracy in floating-point arithmetic. The modification involves exploiting the values $\|A^{k}\|^{\frac{1}{k}}$ and the improvements are more evident for upper triangular matrices.
The (currently) most popular method to compute the $\varphi$-functions was proposed by the same authors in 2011 [al2011computing]. In fact, the purpose of that paper is twofold. The authors observe first that the problem of computing sums of the form $\sum_{k=0}^{p}\varphi_{k}(A)u_{k}$ functions defining an exponential method can be recasted as a single matrix exponential $e^{\tilde{A}}$ where $\tilde{A}$ is obtained by suitably augmenting $A$ with $p$ rows and columns. Next, they propose their algorithm for computing products of matrix exponential times matrices $\mathrm{e}^{A}B$ for $B\in\mathbb{R}^{n\times m}$ with $m\ll n$, as well as $\mathrm{e}^{tA}B$ where $t$ is an array of equally-spaced time points. It is based on the scaling part of the scaling and squaring, plus a truncated Taylor series.
@@
Optimal values for $s$ and $m$ are obtained through a backward error analysis. The algorithm incorporates the ideas previously proposed in [al2010new] in order to minimize overscaling and turns out to be generally superior to existing ones in terms of accuracy, stability and predictability of cost, since the only convergence test involved is a test for early termination built into the evaluation of the truncated Taylor series. [^3] We remark that the algorithm is implemented in the MATLAB function expmv.
@@
$u_{1}$ $=(4.54\times 10^{-5})\times 1+0.01\varphi_{1}(-10)\times 1\approx 5.54\times 10^{-5}\,.$
Now, we can compare the performance of ETD Euler and the classical order 4 RK method (RK4), whose application to (6) gives:
@@
Observe that the origin is a stable steady state of a numerical method for the linear equation whenever $|h\lambda|$ lies in the corresponding stability region. Thus, in general, stability of the origin depends on the choice of $h$. A desirable property of a numerical method is $A$-stability: a scheme is $A$-stable [Dahlquist1963Astab] if the entire negative complex half plane is contained in its stability domain, such that the stability of the origin is independent of the chosen step size.